OpenStudy (anonymous):

Assume that y varies inversely with x and that y is 2 when x is 4. Which of the following is the equation of inverse variation? A) y = 2x B) xy = 8 C) y/x = 2 Fill in the blank. If xy = 20, then y decreases as x _________. A) decreases B) increases C) stays constant Which equation is the inverse variation relationship for the following quantities? y =-12 when x = 5 A) y/x = -2.4 B) y = -2.4x C) xy = -60 Assume that y varies inversely with x. If y = -3 when x = 2.4, what is y when x = 2? A) -3.6 B) -0.8 C) -7.2 D) 4.8 These are the questions i dont get on my practice exam....anyone help???

OpenStudy (anonymous):

the answer have to select from the list?

OpenStudy (anonymous):

\[\frac{2}{y}=\frac{x}{4}\] cross and mutiply xy=8 is B)

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

sencond: xy=20 y =20-x answer is : A)

OpenStudy (anonymous):

third:\[\frac{-12}{y}=\frac{x}{5}\] cross mutipfly xy= -60 is: C)

OpenStudy (anonymous):

You dont have to do all of theem

OpenStudy (anonymous):

last one\[\frac{-3}{y}=\frac{2}{2.4}\] cross time \[y=\frac{-3\times2.4}{2}=-3.6\]

OpenStudy (anonymous):

thank ,but late

OpenStudy (anonymous):

Thank You soo muchh!!!!!!